# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v4

@@ 6 Evaluation of Agentic AI Systems
Trustworthiness. This criterion evaluates reliability, safety, and ethical alignment [159]. Core signals include task success rates under distributional shifts (robustness), violation rate of safety/ethical rules (lower is better), and calibration quality (e.g., Expected Calibration Error, ECE). To avoid pathological behavior in scoring, we define a normalized trustworthiness index
$T\;=\;\frac{w_{C}\,C+w_{R}\,R+w_{I}\,I}{\,1+\lambda\,S\,}\,$, (1)
where $C$ (credibility/accuracy), $R$ (reliability over time), and $I$ (user alignment/rapport) are normalized to $[0,1]$, $S\!\geq\!0$ is self-orientation (higher is worse), $w_{C},w_{R},w_{I}\!\geq\!0$ with $w_{C}{+}w_{R}{+}w_{I}\!=\!1$, and $\lambda\!\geq\!0$ controls how strongly self-orientation penalizes trust.
@@
Explainability. Explainability metrics assess how well human evaluators can understand and trace decisions. Useful measures include coverage (fraction of outputs with an explanation), fidelity/faithfulness (agreement with the model’s actual decision basis), stability (consistency across similar cases), and human interpretability ratings. Suites like OpenXAI [160] report faithfulness, stability, and fairness of explanations. High explainability builds user trust and aids debugging; in regulated domains (e.g., healthcare, finance) it is often essential for compliance.
@@
User-Centered Performance. User-centered metrics capture how effectively the agent satisfies end-user needs [161]. Signals include post-interaction satisfaction scores, goal-fulfillment rate, number of clarification turns (lower is better), and perceived coherence/naturalness. Human-in-the-loop (HITL) studies typically rate helpfulness, clarity, and instruction adherence. Ultimately, a user-centered agent aligns actions with user intent and preferences.
Table 5: Summary of key evaluation metrics for agentic AI systems. Each category addresses a distinct facet of performance and behavior.
Metric Aspect Evaluation Focus (Examples of Metrics)
Trustworthiness Reliability, safety, and alignment. Examples: success rate under shifts; safety/ethics violation rate (lower is better); calibration (ECE/Brier  [^9] https://en.wikipedia.org/wiki/Brier_score); fairness indices; red-team/jailbreak success (lower is better).
Explainability Transparency and traceability. Examples: explanation coverage; faithfulness/fidelity; stability across similar cases; human interpretability scores; explanation time/cost.
User-Centered Experience and outcomes. Examples: satisfaction (Likert/CSAT  [^10] https://www.ibm.com/think/topics/csat-customer-satisfaction-score); goal-completion rate; clarification turns; perceived coherence/naturalness; task time; cognitive load (NASA-TLX).
Coordination Collaboration in multi-agent or modular setups. Examples: team task completion; coordination overhead (messages/tokens, rounds to converge); plan consistency; deadlock/conflict rate (lower is better); synergy score for complementary actions.
Composite Holistic summaries. Examples: weighted aggregate across categories (weights sum to 1); Tool-Use Efficacy (TUE, see definition); cross-domain benchmark aggregates.
@@
Coordination. In multi-agent or modular systems, coordination captures whether components maintain a shared plan, minimize redundant work, and resolve dependencies. Quantify with team success rate, communication efficiency (messages/tokens and rounds to consensus), plan/belief consistency, deadlock/conflict rate, and a Component Synergy Score (CSS) that credits inter-agent actions which enable or improve peers’ outcomes (higher is better).
Composite Metrics.
@@
$M_{\text{composite}}\;=\;w_{T}M_{T}+w_{E}M_{E}+w_{U}M_{U}+w_{C}M_{C}\,$, (2)
where $M_{T},M_{E},M_{U},M_{C}\in[0,1]$ are normalized category scores for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}\!\geq\!0$ with $\sum w_{\bullet}=1$ encode domain priorities. While this provides a scalar index for ease of interpretation, it risks masking over-performance in one dimension with under-performance in another. To address this, we treat the evaluation more rigorously as a vector of metrics,
$\mathbf{M}\;=\;(M_{T},M_{E},M_{U},M_{C})\,$, (3)
and recommend selection based on $\arg\max$ over $\mathbf{M}$ rather than on any scalarization:
$\arg\max_{i}\;\mathbf{M}_{i}\,.$ (4)
In this view, Eq. 2 can be interpreted as a limiting case, useful for communicating results to non-technical stakeholders, while Eq. 3 and Eq. 4 form the principled basis for evaluation and comparison.
Definition (Tool-Use Efficacy, TUE). In this work, we introduce Tool-Use Efficacy (TUE) as a specialized composite for evaluating how correctly and efficiently an agent uses external tools. Let
[…]
$\mathrm{Exec}$ $\in[0,1]$ $\text{(execution success: no exceptions/timeouts, correct API usage)},$
$\mathrm{Out}$ $\in[0,1]$ $\text{(post-call outcome correctness on tool-gated sub-tasks)},$
$\mathrm{Eff}$ $\in[0,1]$ $\text{(efficiency: normalized inverse cost for latency/tokens/\#calls)}.$
We define two aggregators (both reportable; choose per application):
(a) Linear aggregator (auditable weights).
$\mathrm{TUE}_{\mathrm{lin}}\;=\;\alpha_{\!s}\,\mathrm{Sel}+\alpha_{\!a}\,\mathrm{Arg}+\alpha_{\!e}\,\mathrm{Exec}+\alpha_{\!o}\,\mathrm{Out}+\alpha_{\!f}\,\mathrm{Eff}\,,\quad\alpha_{\!*}\!\geq\!0,\;\sum\alpha_{\!*}=1.$ (5)
(b) Geometric aggregator (“all-of-the-above” strictness).
$\mathrm{TUE}_{\mathrm{geo}}\;=\;\Big{(}\mathrm{Sel}^{\beta_{s}}\,\mathrm{Arg}^{\beta_{a}}\,\mathrm{Exec}^{\beta_{e}}\,\mathrm{Out}^{\beta_{o}}\,\mathrm{Eff}^{\beta_{f}}\Big{)}^{\!1/(\beta_{s}{+}\beta_{a}{+}\beta_{e}{+}\beta_{o}{+}\beta_{f})},\quad\beta_{\!*}\!>\!0.$ (6)
Overhead penalty. If the preference is a separate cost term, then we can define normalized overhead $\mathrm{OH}\in[0,1]$ (increasing in time/tokens/calls) and report the penalized score
$\mathrm{TUE}^{\star}\;=\;\frac{\mathrm{TUE}}{\,1+\lambda\,\mathrm{OH}\,}\,,\quad\lambda\!\geq\!0,$ (7)
where $\mathrm{TUE}$ is either (5) or (6). It is recommended to always report the components (Sel, Arg, Exec, Out, Eff) with confidence intervals to ensure reproducibility.
@@ 7 Security and Privacy in Agentic AI Systems
In this section, we discuss the security and privacy in AMAS.
@@ 7.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [162]. Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [163]. Among the foundational techniques employed are encryption , access control , adversarial defense , and runtime monitoring, each adapted to the unique demands of decentralized multi-agent environments.
Encryption. Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [67]. Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption , and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
Access control. Access conrol is highly important when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [57]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [157] and Attribute-Based Access Control (ABAC) [157] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
Adversarial learning. Adversarial learning is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results. Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially designing outputs from one compromised agent propagating misleading information to others [164]. Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities. Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
Runtime Monitoring. Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity. Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [165]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
@@ 7.2 Privacy-Preserving Techniques
Differential privacy (DP). DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [166]. DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev [30] or ReAct-based pipelines [75] are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [167] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC) [168], homomorphic encryption [169], and trusted execution environments (TEEs) [^11] Trusted Execution Environment (Wikipedia) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
Privacy-by-design [170] principles are becoming central to the engineering of next-generation Agentic AI systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
Next, we discuss the compliance to AI goverance and policy in Agentic AI systems.
@@
Governance Org. accountability Define owners for models/agents/tools; RACI for changes; approvals for high-impact actions Governance policy; RACI records; change approvals NIST AI RMF (Govern) [171]; ISO/IEC 42001 [2]
Governance Policy-as-code Encode allow/deny rules in orchestrator; pre-exec checks and human sign-off gates Policy repository; policy test cases; gate logs EU AI Act Arts. 11–14 [125]; ISO/IEC 42001 [2]
Risk Mgmt Risk register & AIA Hazard/threat enumeration; likelihood/impact scoring; AI Impact Assessment (AIA) Risk register; AIA reports; mitigation plan ISO/IEC 23894; ISO/IEC 42005 [129]; EU AI Act Art. 9 [125]
Risk Mgmt Post-market monitoring Drift/jailbreak telemetry; incident thresholds; rollback playbooks PMM plan; incident tickets; root-cause analysis EU AI Act Arts. 72–73 [125]; NIST AI RMF (Manage) [171]
Data Gov Data minimization & lineage Task/tenant-scoped memory; TTLs/redaction; dataset lineage and licenses Data inventory; lineage graphs; retention logs EU GDPR Art. 25 [34]; ISO/IEC 5259 series
Data Gov Quality & bias checks Coverage/balance checks; label noise audits; sampling documentation Data quality reports; bias audit results EU AI Act Art. 10 [125]; ISO/IEC TR 24027
Explainability Faithfulness/stability Log rationales; test explanation faithfulness/stability; user-facing summaries XAI eval report; model/data/agent cards NISTIR 8312 [155]; NIST AI RMF (Map/Measure) [171]
Documentation Technical documentation Annex-IV tech docs: purpose, data, performance, oversight, foreseeable misuse Annex-IV dossier; limitations & misuse notes EU AI Act Arts. 11–12; Annex IV [125]
Traceability Provenance & logging Log prompts, plans, tool calls, I/O, timestamps, agent role; sign releases Immutable logs (WORM/tamper-evident); signed artifacts EU AI Act Arts. 12–13 [125]; ISO/IEC 42001 [2]
Human Oversight Human-in-the-loop gates Require human approval for sensitive plans (delete, external write, PII) Approval records; override/rollback logs EU AI Act Art. 14 [125]; NIST AI RMF (Manage) [171]
Robustness Stress/red-team testing Adversarial prompts; perturbation/fault tests; safety regression suites Red-team report; robustness curves; residual risk EU AI Act Art. 15 [125]; ISO/IEC 24029-1 [32]
Security I/O mediation Prompt sanitization; output filtering; sensitive-data (PII/PHI/secrets) detection Filter policies; violation/egress logs OWASP LLM Top-10 [137]; HIPAA §164 [35]
@@
Privacy DP budgeting & consent Differential privacy budgets; consent bases; data-subject rights flows DP budget ledger; DSR tickets; consent records GDPR Art. 25 [34]; OECD principles [127]
@@
Coordination Role separation & least privilege Distinct agent roles (planner, coder, reviewer); minimal cross-role permissions Role definitions; privilege matrices; audit samples NIST AI RMF (Govern) [171]; ISO/IEC 42001
Coordination Plan consistency checks Shared plan/belief verification; conflict/deadlock detection; CSS metric Plan snapshots; CSS scores; conflict logs Multi-agent eval practice; robustness [32]
Monitoring Anomaly & exfil detection Detect unusual tool chains, volume spikes, strange domains; rate-limit/kill switch Anomaly dashboards; blocked-egress records OWASP LLM Top-10 [137]; NIST AI RMF (Manage)
Transparency User notices & recourse User-facing disclosures, capabilities/limits, and appeal/feedback channels Notices; feedback logs; SLA for appeals OECD principles [127]; EU AI Act transparency
Evaluation Fairness & calibration Group metrics; ECE/Brier; CI reporting; risk–benefit trade-off docs Fairness report; calibration plots; CI tables NIST AI RMF (Measure) [171]; ISO/IEC TR 24027
Records Release/rollback discipline Versioned releases; rollback points; emergency disable/recall Release notes; rollback attestations EU AI Act PMM [125]; ISO/IEC 42001
GPAI GPAI obligations Provide model cards, training-data summary, eval reports, and usage restrictions; disclose known limitations GPAI documentation pack; model card; eval report EU AI Act (GPAI duties) [125]; ISO/IEC 42001
Transparency Synthetic content disclosure Label AI-generated/edited media; attach provenance (e.g., C2PA) and user notices Watermarking/provenance config; disclosure logs EU AI Act Art. 52 [125]; C2PA121212c2pa.org
Privacy Cross-border transfers SCCs/TIAs for third-country transfers; data-residency controls; key management SCC/TIA records; DPA; data-map; key-custody logs GDPR Ch. V [34]
Privacy Data-subject rights ops Verified workflows for access, erasure, rectification, restriction, portability, objection DSR tickets/SLAs; fulfillment proofs; audit samples GDPR Arts. 12–23 [34]
Operations Change mgmt & release gates Risk-based pre-release checklist; dual-control approvals; staged rollouts; rollback criteria Change tickets; gate results; approvals; rollback points ISO/IEC 42001; NIST AI RMF (Manage) [171]
@@
Human Factors HCD & accessibility Human-centred design reviews; usability risk analysis; accessibility conformance checks HCD review notes; usability test reports; a11y checklists ISO 9241-210131313ISO 9241-210; OECD principles [127]

### v5

@@ 5 TRiSM Evaluation of Agentic AI Systems
This dimension evaluates reliability, safety, and ethical alignment [liu2023gevalnlgevaluationusing]. Core signals include success under distributional shift (robustness), safety/ethical violation rate (lower is better), and calibration quality (e.g., ECE/Brier). To avoid pathological scoring, we define a normalized trustworthiness index:
$T\;=\;\frac{w_{\text{acc}}\,A+w_{\text{rob}}\,R+w_{\text{align}}\,L}{1+\lambda\,V}\,$, (1)
where $A$ is calibrated task success (accuracy adjusted for calibration), $R$ is robustness under shifts, $L$ is alignment/safety compliance (e.g., 1violation rate1{-$\text${violation rate}), all normalized to $[0,1]$, and $V\!\geq\!0$ is a normalized violation/self-serving penalty term (higher is worse). The weights satisfy $w_{\text{acc}}{+}w_{\text{rob}}{+}w_{\text{align}}=1$ and $\lambda\!\geq\!0$ controls how strongly violations penalize trust.
Explainability
@@
Explainability metrics assess how well humans can understand and trace decisions. Useful measures include coverage (fraction of outputs with explanations), faithfulness/fidelity (agreement with the model’s actual decision basis), stability (consistency under small perturbations), and human interpretability ratings. Suites such as OpenXAI [agarwal2024openxaitransparentevaluationmodel] report faithfulness, stability, and fairness-related properties of explanations. High explainability supports debugging and, in regulated domains (e.g., healthcare and finance), is often necessary for compliance.
User-Centered Performance
@@
User-centered metrics capture how effectively the agent satisfies end-user needs [rodden2010measuring]. Signals include post-interaction satisfaction (e.g., CSAT), goal-fulfillment rate, number of clarification turns (lower is better), and perceived coherence/naturalness. Human-in-the-loop (HITL) studies commonly rate helpfulness, clarity, and instruction adherence. A user-centered agent aligns actions with user intent and preferences.
Table 6: Summary of key evaluation dimensions for Agentic AI systems.
Aspect Focus Strengths Limitations
Trustworthiness Reliability, safety, alignment (OOD success, safety violations, calibration, fairness). Deployment readiness; regulatory alignment (EU AI Act); statistically grounded; captures safety failures. Assumes stationarity; fairness trade-offs across groups [raza_fair_2024]; incomplete edge-case coverage.
Explainability Transparency and traceability (coverage, fidelity, stability, human interpretability). Supports debugging and compliance [ai2023artificial]; improves calibrated trust; standardized benchmarks (OpenXAI [agarwal2024openxaitransparentevaluationmodel]). Expensive fidelity checks; faithful $\neq$ interpretable; weak multi-agent attribution.
User-Centered User experience and outcomes (CSAT, task success, clarification turns, NASA-TLX). Reflects real-world utility; identifies usability bottlenecks; validated human factors measures. Subjective and high variance; costly user studies; limited scalability.
Coordination Multi-agent collaboration (task success, communication cost, plan consistency, deadlocks). Captures emergent AMAS behavior; measures coordination efficiency; decomposable contribution metrics. No standard benchmarks; sensitive to orchestration; may penalize useful redundancy.
Composite Aggregate system performance (weighted scores, tool-use efficacy). Single actionable summary; enables cross-system comparison; supports thresholding. Subjective weighting; can mask single-dimension failures; prone to Goodhart effects.
Coordination
@@
In multi-agent or modular systems, coordination measures whether components maintain a shared plan, minimize redundant work, and resolve dependencies. Common signals include team success rate, communication efficiency (messages/tokens and rounds-to-consensus), plan/belief consistency, and deadlock/conflict rate. We additionally report: (i) a Component Synergy Score (CSS) to credit enabling contributions across agents, and (ii) Tool Utilization Efficacy (TUE) to assess correctness and efficiency of tool calls.
Composite Reporting
@@
$M_{\text{comp}}\;=\;\omega_{T}M_{T}+\omega_{E}M_{E}+\omega_{U}M_{U}+\omega_{Co}M_{Co}\,$, (2)
where $M_{T},M_{E},M_{U},M_{Co}\in[0,1]$ are normalized scores for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $\omega_{\bullet}\geq 0$ with $\sum\omega_{\bullet}=1$ encode domain priorities. Since scalarization can hide critical failures, we recommend always reporting the full metric vector:
$\mathbf{M}\;=\;(M_{T},M_{E},M_{U},M_{Co})\,.$ (3)
Model selection should then be based on (i) minimum thresholds (e.g., $M_{T}\geq\tau_{T}$ and $M_{Co}\geq\tau_{Co}$) and/or (ii) Pareto-optimality (non-dominated models) rather than relying only on a single weighted score.
Component Synergy Score (CSS)
Definition: CSS quantifies the quality of inter-agent collaboration by measuring how effectively one agent’s actions enable or enhance a peer’s downstream performance.
Formal expression:
$\text{CSS}=\frac{1}{N_{\text{pairs}}}\sum_{i}\sum_{j\neq i}\Big[\text{Impact}(a_{i}\!\rightarrow\!a_{j})\cdot\text{Quality}(a_{j}\mid a_{i})\Big],$ (4)
where $N_{\text{pairs}}$ is the number of ordered agent pairs $(i,j)$ included in the evaluation. $\text{Impact}(a_{i}\!\rightarrow\!a_{j})$ measures how $a_{i}$ changes $a_{j}$’s subsequent performance (e.g., relative improvement), and $\text{Quality}(a_{j}\mid a_{i})$ is the quality of $a_{j}$’s output conditioned on $a_{i}$’s contribution.
Illustrative example: In a software AMAS with Planner, Coder, and Tester:
  • Planner’s decomposition speeds up Coder by 35% ($\text{Impact}=0.35$).
  • Tester achieves 92% coverage conditioned on Coder’s output ($\text{Quality}=0.92$).
If we score only this link, then $\text{CSS}=0.35\times 0.92\approx 0.32$ (and increases as additional enabling links are included).
Tool Utilization Efficacy (TUE)
Definition: TUE evaluates how correctly and efficiently agents invoke external tools (APIs, databases, code interpreters) within their workflows.
Components:
  • $\text{Sel}\in[0,1]$ — Selection (decision precision/recall for tool use)
  • $\text{Arg}\in[0,1]$ — Argument validity (well-formed parameters)
  • $\text{Exec}\in[0,1]$ — Execution success (no errors/timeouts)
  • $\text{Out}\in[0,1]$ — Outcome correctness (downstream task correctness)
  • $\text{Eff}\in[0,1]$ — Efficiency (normalized inverse of latency/token/call cost)
Linear aggregator:
$\text{TUE}_{\text{lin}}=\alpha_{s}\text{Sel}+\alpha_{a}\text{Arg}+\alpha_{e}\text{Exec}+\alpha_{o}\text{Out}+\alpha_{f}\text{Eff},$ (5)
where $\alpha_{\bullet}\geq 0$ and $\sum\alpha_{\bullet}=1$.
Key Insights
This section defined a multi-dimensional evaluation framework covering trustworthiness (robustness, violation rates, calibration), explainability (coverage, faithfulness, stability), user-centered performance (satisfaction and goal fulfillment), and coordination quality. CSS captures positive inter-agent enablement, while TUE decomposes tool use into selection, argument validity, execution success, outcome correctness, and efficiency. Finally, the framework emphasizes reporting a metric vector and using thresholding/Pareto analysis to reduce the risk that scalar composites mask critical single-dimension failures.
@@ 6 Security and Privacy in Agentic AI Systems
Evaluating the security and robustness of AMAS requires standardized benchmarks and well-defined metrics to assess both attack vulnerabilities and defense effectiveness. This subsection reviews current evaluation frameworks and security metrics for agentic AI, aligned with the TRiSM pillars.
@@ 6.1 Security Mechanisms
Agentic AI systems (multiple collaborating LLM agents) have a bigger attack surface than traditional agents, so they need layered defenses: protecting data, execution, communication, and model robustness; using encryption, access control, adversarial defenses, and runtime monitoring tailored to multi-agent setups.
Encryption. Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [feretzakis2024privacy]. Agentic workflows often include the inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption , and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
Access control. Access control is highly important when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [wu2023autogen]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [khan2024role] and Attribute-Based Access Control (ABAC) [khan2024role] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
Adversarial learning. Adversarial threats are increasing for LLM-based agents, including prompt injection, poisoned tool outputs, and disruption via malformed intermediate results. Recent work shows that compromising one agent can propagate misleading outputs and destabilize an entire multi-agent framework [zhang2024breaking]. Mitigations include adversarial training (e.g., input perturbation, reward shaping, contrastive learning) and practical safeguards such as enforcing safety constraints and validating tool outputs before execution.
Runtime Monitoring. Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity. Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [zhen2025novel]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
@@ 6.2 Privacy-Preserving Techniques
Differential privacy (DP). DP protects individuals by adding controlled noise, and in multi-agent LLM systems it can be used during training or inference to limit what agents can reveal to each other. Using approaches like DP in federated learning, DP-SGD, and an $\epsilon$ privacy budget, agents can collaboratively train, fine-tune, or coordinate while reducing exposure of raw or identifiable data [lazaros2024federated].
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev [qian2023chatdev] or ReAct-based pipelines [yao2023reactsynergizingreasoningacting] are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [luzon2024tutorial] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC) [evans2018pragmatic], homomorphic encryption [acar2018survey], and trusted execution environments (TEEs) [^13] Trusted Execution Environment (Wikipedia) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
Privacy-by-design [schaar2010privacy] principles are becoming central to the engineering of next-generation Agentic AI systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
@@
Governance Org. accountability Define owners for models/agents/tools; RACI for changes; approvals for high-impact actions Governance policy; RACI records; change approvals NIST AI RMF (Govern) [tabassi2023ai]; ISO/IEC 42001 [iso42001]
Governance Policy-as-code Encode allow/deny rules in orchestrator; pre-exec checks and human sign-off gates Policy repository; policy test cases; gate logs EU AI Act Arts. 11–14 [eu_ai_act]; ISO/IEC 42001 [iso42001]
Risk Mgmt Risk register & AIA Hazard/threat enumeration; likelihood/impact scoring; AI Impact Assessment (AIA) Risk register; AIA reports; mitigation plan ISO/IEC 23894; ISO/IEC 42005 [ISO_IEC_42005_2025]; EU AI Act Art. 9 [eu_ai_act]
Risk Mgmt Post-market monitoring Drift/jailbreak telemetry; incident thresholds; rollback playbooks PMM plan; incident tickets; root-cause analysis EU AI Act Arts. 72–73 [eu_ai_act]; NIST AI RMF (Manage) [tabassi2023ai]
Data Gov Data minimization & lineage Task/tenant-scoped memory; TTLs/redaction; dataset lineage and licenses Data inventory; lineage graphs; retention logs EU GDPR Art. 25 [gdpr25]; ISO/IEC 5259 series
Data Gov Quality & bias checks Coverage/balance checks; label noise audits; sampling documentation Data quality reports; bias audit results EU AI Act Art. 10 [eu_ai_act]; ISO/IEC TR 24027
Explainability Faithfulness/stability Log rationales; test explanation faithfulness/stability; user-facing summaries XAI eval report; model/data/agent cards NISTIR 8312 [phillips2021four]; NIST AI RMF (Map/Measure) [tabassi2023ai]
Documentation Technical documentation Annex-IV tech docs: purpose, data, performance, oversight, foreseeable misuse Annex-IV dossier; limitations & misuse notes EU AI Act Arts. 11–12; Annex IV [eu_ai_act]
Traceability Provenance & logging Log prompts, plans, tool calls, I/O, timestamps, agent role; sign releases Immutable logs (WORM/tamper-evident); signed artifacts EU AI Act Arts. 12–13 [eu_ai_act]; ISO/IEC 42001 [iso42001]
Human Oversight Human-in-the-loop gates Require human approval for sensitive plans (delete, external write, PII) Approval records; override/rollback logs EU AI Act Art. 14 [eu_ai_act]; NIST AI RMF (Manage) [tabassi2023ai]
Robustness Stress/red-team testing Adversarial prompts; perturbation/fault tests; safety regression suites Red-team report; robustness curves; residual risk EU AI Act Art. 15 [eu_ai_act]; ISO/IEC 24029-1 [iso24029-1]
Security I/O mediation Prompt sanitization; output filtering; sensitive-data (PII/PHI/secrets) detection Filter policies; violation/egress logs OWASP LLM Top-10 [owasp2024llmtop10]; HIPAA §164 [hipaa164]
@@
Privacy DP budgeting & consent Differential privacy budgets; consent bases; data-subject rights flows DP budget ledger; DSR tickets; consent records GDPR Art. 25 [gdpr25]; OECD principles [oecd_ai_catalogue]
@@
Coordination Role separation & least privilege Distinct agent roles (planner, coder, reviewer); minimal cross-role permissions Role definitions; privilege matrices; audit samples NIST AI RMF (Govern) [tabassi2023ai]; ISO/IEC 42001
Coordination Plan consistency checks Shared plan/belief verification; conflict/deadlock detection; CSS metric Plan snapshots; CSS scores; conflict logs Multi-agent eval practice; robustness [iso24029-1]
Monitoring Anomaly & exfil detection Detect unusual tool chains, volume spikes, strange domains; rate-limit/kill switch Anomaly dashboards; blocked-egress records OWASP LLM Top-10 [owasp2024llmtop10]; NIST AI RMF (Manage)
Transparency User notices & recourse User-facing disclosures, capabilities/limits, and appeal/feedback channels Notices; feedback logs; SLA for appeals OECD principles [oecd_ai_catalogue]; EU AI Act transparency
Evaluation Fairness & calibration Group metrics; ECE/Brier; CI reporting; risk–benefit trade-off docs Fairness report; calibration plots; CI tables NIST AI RMF (Measure) [tabassi2023ai]; ISO/IEC TR 24027
Records Release/rollback discipline Versioned releases; rollback points; emergency disable/recall Release notes; rollback attestations EU AI Act PMM [eu_ai_act]; ISO/IEC 42001
GPAI GPAI obligations Provide model cards, training-data summary, eval reports, and usage restrictions; disclose known limitations GPAI documentation pack; model card; eval report EU AI Act (GPAI duties) [eu_ai_act]; ISO/IEC 42001
Transparency Synthetic content disclosure Label AI-generated/edited media; attach provenance (e.g., C2PA) and user notices Watermarking/provenance config; disclosure logs EU AI Act Art. 52 [eu_ai_act]; C2PA141414c2pa.org
Privacy Cross-border transfers SCCs/TIAs for third-country transfers; data-residency controls; key management SCC/TIA records; DPA; data-map; key-custody logs GDPR Ch. V [gdpr25]
Privacy Data-subject rights ops Verified workflows for access, erasure, rectification, restriction, portability, objection DSR tickets/SLAs; fulfillment proofs; audit samples GDPR Arts. 12–23 [gdpr25]
Operations Change mgmt & release gates Risk-based pre-release checklist; dual-control approvals; staged rollouts; rollback criteria Change tickets; gate results; approvals; rollback points ISO/IEC 42001; NIST AI RMF (Manage) [tabassi2023ai]
@@
Human Factors HCD & accessibility Human-centred design reviews; usability risk analysis; accessibility conformance checks HCD review notes; usability test reports; a11y checklists ISO 9241-210151515ISO 9241-210; OECD principles [oecd_ai_catalogue]